# Customizing Text-to-Image Models with a Single Image Pair

### v1

@@ 4.3 Results
Quantitative evaluation.   We show quantitative results in Figure 4. Increased marker size (circles) indicates the higher application of style, and line color determines the method. When evaluating style similarity vs. structure preservation in Figure 4, we see that our training method’s Pareto dominates all baselines, yielding lower perceptual distance to style images while still being perceptually similar to the original content image. Secondly, style guidance outperforms the LoRA scale for Ours (blue vs orange), DB LoRA (green vs. pink), and Concept Sliders (brown vs. purple), highlighting the effectiveness of both parts of our method.
@@
Qualitative evaluation.   We compare our method with the highest performing baselines in Figure 6. The finetuning-based methods DB LoRA [33, 74] and Concept Sliders  [22] outperform the encoder-based method [95] for our task. Hence, we compare against that in Figure 6. For both baselines, we modulate style application with LoRA scale (Equation 3). We observe that DB LoRA often fails to generate the style-transformed version of the original image and overfits to the training pair image when generating similar concepts. There are two main reasons why this may occur. First, we are in a challenging case where there is only $1$ training image instead of the usual $3-5$ images that customization methods use. Second, we are prompting the model on the same or very similar text prompts to the training prompt, and the baseline method overfits to the training image for these prompts. Our method preserves the structure of the original image while applying the learned style. Moreover, applying our style guidance instead of the LoRA scale benefits the baseline method as well (Figure 6, last $2$ columns), as it can better preserve the structure of the original image, though it still tends to overfit to the content of the training image. We observe a similar issue for other baselines as well. We show a qualitative comparison with the other baselines in Appendix 13. We also compare with baselines using our style guidance for style application at inference time in Appendix 13.
@@
User preference study.   We perform a user preference study using Amazon Mechanical Turk. We test our method against all baselines, as well as a version of our method trained without orthogonality constraint. Specifically, we test on all datasets in Section 4.1. When evaluating against DB LoRA and Concept Sliders, we consider inference with both LoRA scale as in Equation 3 and style guidance as in Equation 10. For each method, we pick a single style strength that performs most optimally according to quantitative metrics as in Figure 4. Full details are available in Appendix 0.C. We collect $400$ responses per paired test of ours vs the other method. The user is shown an image generated via our method and an image generated via the other method and asked to select the image that best applies the given style to the new content image. We provide a detailed setup of the user study in Appendix 0.C. As shown in Figure 5, our method is favored by users in comparison to baselines, whether evaluating images generated within the same category as the training image pair or across different categories. Secondly, users prefer our full method to ours without the orthogonality constraint.
@@
Figure 7: Blending multiple style guidances. We can compose multiple customized models by directly blending each style guidance together. Adjusting blending strength of each model allows us to acquire a smooth style transition. Train Image Pair 0 Style image credits: Jack Parkhouse
@@
Figure 8: Limitations. Left: our method struggles with categories when they significantly differ from the training categories. Here, our method fails to transfer the artistic style of landscape image pairs to human portraits. Right: our method can cause structure changes in some instances, like change of body position or background changes.
@@ 5 Discussion and Limitations
Limitations.   First, while our method is able to transfer the style from a pair of dog images to cat photos, it struggles to handle completely different categories from the training image pair, particularly when the test category significantly differs from the training. As shown in Figure 8 (left), our method falls short of faithfully replicating the style of landscape paintings in some human images.
@@
Finally, our method may occasionally fail to completely maintain input structure, as demonstrated in Figure 8 (right).
Acknowledgments.   We would like to thank Ali Jahanian, Gaurav Parmar, Ruihan Gao, Sean Liu, and Or Patashnik for their insightful feedback and input that contributed to the finished work. We also thank Jack Parkhouse and Aaron Hertzmann for providing style images. Maxwell Jones is supported by the Siebel Scholar fellowship. This project is partly supported by the Amazon Faculty Research Award, NSF IIS-2239076, Open Philanthropy, and the Packard Fellowship.
@@
In Section 0.A, we evaluate our method against baselines on the diversity metric, showing that our method leads to more diverse generations comparatively. We also show more qualitative results along with a comparison to the concurrent work of Style Aligned Image Generation [28]. In Section 0.B, we then present details of our style guidance formulation. Finally, in Section 0.C, we provide more implementation details, including the setup for our human preference study and the full synthetic training dataset used for evaluation.
Appendix 0.A More Quantitative and Qualitative Results
Diversity metric.   To measure the overfitting behavior of our method and baselines, we consider a diversity metric. Concretely, we measure the DreamSim [17] perceptual distance between any two images trained with the same style image pair and generated with the same prompt and average results over training pairs and prompts. More formally, we let
DreamSim $\text{DreamSim Diversity}=\mathbb{E}_{S\in\mathcal{S},P\in\mathcal{P}}\left[\mathbb{E}_{i_{1},i_{2}\in\text{data}_{S,P}}\text{DreamSim}(i_{1},i_{2})\right]$ (12)
@@ Appendix 0.A More Quantitative and Qualitative Results
Figure 12: Result of our method compared to the methods without finetuning (zoom in for best viewing). For all methods, we consider adding the edgemap from the pretrained output as an extra conditioning using ControlNet. Without this edgemap, other methods tend to lose the structure of the pretrained output. In some cases, however, an additional edgemap can overly constrain the output of a model, like in the second and fourth stylistic image pairs. Our method preserves the structure of the Stable Diffusion image, while faithfully applying the desired style. We use style guidance strength 3 and classifier guidance strength 5 for our method and set the IP-adapter scale and style-alignment scale to 0.5.
Style Aligned Image Generation [28] Baseline   This is a recent work for zero-shot style-consistent image generation from an exemplar style image. Given the exemplar style image, it is first inverted to a noise map; then for a new text prompt, the image is generated by attending to both its own self-attention map and the self-attention map from the style exemplar at every denoising step. We compare against this baseline by using the style image in our training image pair as an exemplar and generating a new style image with a new text prompt using this method. Optionally, we condition this generation on the edge map of the newly generated image without attention sharing using ControlNet [97] to help with content preservation. We show the qualitative results of our method compared to all the variants of this baseline in Figure 12. Figures 10 and 11 show quantitative comparison, where our method outperforms this baseline in terms of both style similarity and diversity metric. We achieve lower perceptual distance to the style ground-truth images, low perceptual distance from content images, and high diversity.
Extra Qualitative Evaluation   We compare our method to non finetuning-based methods in Figure 12. We observe that these methods perform worse than finetuning-based methods, especially when generating images in a different category to the training style image. Secondly, we compare our method with the highest-performing baselines, but use our style guidance (Equation 10 ) to apply stylization during inference for these baselines. We present our results in Figure 13. First, we notice that using style guidance for adding style allows the baseline methods to better preserve original content over LoRA scale (Figure 13 vs Figure 6). While adding our style guidance is better able to preserve content while applying style for baseline methods, our full method is still able to outperform baselines with style guidance applied.
Appendix 0.B Style Guidance Details
@@
$P({\mathbf{x}}|c_{t},c_{\text{style}})=\frac{P({\mathbf{x}},c_{t},c_{\text{style}})}{P(c_{t},c_{\text{style}})}=\frac{P(c_{\text{style}}|c_{t},{\mathbf{x}})P(c_{t}|{\mathbf{x}})P({\mathbf{x}})}{P(c_{t},c_{\text{style}})}$ (13)
@@
$=$ $\ \nabla_{\mathbf{x}}\log\left(\frac{P(c_{\text{style}},c_{t},{\mathbf{x}})}{P(c_{t},{\mathbf{x}})}\right)+\nabla_{\mathbf{x}}\log\left(\frac{P(c_{t},{\mathbf{x}})}{P({\mathbf{x}})}\right)+\nabla_{\mathbf{x}}\log\left(P({\mathbf{x}})\right)$
@@
[…]
$\hat{\epsilon}_{\theta}({\mathbf{x}}_{t},c_{t},{\mathbf{c}}_{\text{style}})=$ $\epsilon_{\theta}({\mathbf{x}}_{t},\varnothing)$
@@
$\lambda_{\text{cfg}}$ and $\lambda_{\text{style}}$ are guidance scales that can be varied as in classifier free guidance [31]. Given a fixed $\lambda_{\text{cfg}}$, we can vary the $\lambda_{\text{style}}$ term as desired to generate an original guidance $\lambda_{\text{cfg}}$ image with varying amounts of style. Notice that at $\lambda_{\text{cfg}}=\lambda_{\text{style}}$, the $\epsilon_{\theta}({\mathbf{x}}_{t},\varnothing)$ terms cancel and we are left with the original classifier guidance.
Appendix 0.C Implementation Details
@@
Table 1: Experiment Hyperparameters. We choose a fixed stylization hyperparameter for our own model and each baseline when generating images for Mechanical Turk. When picking a hyperparameter, we try and optimize tradeoffs between style application and content preservation, informed by Figure 4 in the main body. Our style guidance (Equation 10) generally takes values from 0 to $\lambda_{\text{cfg}}=5$, while all other stylization hyperparameters generally take values 0 to 1.

### v2

@@ 4.3 Results
Quantitative evaluation.   We show quantitative results against the highest performing baselines in Figure 7. Increased marker size (circles) indicates the higher application of style, and line color determines the method. When evaluating style similarity vs. structure preservation in Figure 7, we see that our training method’s Pareto dominates all baselines, yielding lower perceptual distance to style images while still being perceptually similar to the original content image. We find that DB LoRA and Styledrop perform similarly, and report Styledrop results in Appendix A. Finally, we consider our method with LoRA scale during inference and other baselines with our style guidance during inference for ablation, and provide results in Appendix A.
@@
Qualitative evaluation.   We compare our method with the highest performing baselines in Figure 5. The finetuning-based methods DB LoRA [36, 78] and Concept Sliders  [23] outperform the encoder-based method [100] for our task. Hence, we compare against that in Figure 5. For both baselines, we modulate style application with LoRA scale (Equation 3). We observe that DB LoRA often fails to generate the style-transformed version of the original image and overfits to the training pair image when generating similar concepts. There are two main reasons why this may occur. First, we are in a challenging case where there is only $1$ training image instead of the usual $3-5$ images that customization methods use. Second, we are prompting the model on the same or very similar text prompts to the training prompt, and the baseline method overfits to the training image for these prompts. Our method preserves the structure of the original image while applying the learned style. Moreover, applying our style guidance instead of the LoRA scale benefits the baseline method as well (Figure 5, last $2$ columns), as it can better preserve the structure of the original image, though it still tends to overfit to the content of the training image. We observe a similar issue for other baselines as well.
We compare our method to non finetuning-based methods in Figure 6. We observe that these methods perform worse than finetuning-based methods, especially when generating images in a different category to the training style image. We also compare with baselines using our style guidance for style application at inference time in Appendix 17 .
@@
User preference study.   We perform a user preference study using Amazon Mechanical Turk. We test our method against all baselines, as well as a version of our method trained without the orthogonality constraint. Specifically, we test on all datasets in Section 4.1. When evaluating against DB LoRA and Concept Sliders, we consider inference with both LoRA scale as in Equation 3 and style guidance as in Equation 10. For each method, we pick a single style strength that performs most optimally according to quantitative metrics as in Figure 7. Full details are available in Appendix D, and a user study with all baselines is available in Appendix 17 . We collect $400$ responses per paired test of ours vs the other method. The user is shown an image generated via our method and an image generated via the other method and asked to select the image that best applies the given style to the new content image. We provide a detailed setup of the user study as well as a user study on baseline methods using our style guidance in Appendix D As shown in Figure 8, our method is favored by users in comparison to baselines, whether evaluating images generated within the same category as the training image pair or across different categories. Secondly, users prefer our full method to ours without the orthogonality constraint, specifically when evaluating on the same category as training.
Figure 9: Real image editing. We can edit real images by inverting a real image into a noisy latent code using a diffusion inversion pipeline [87, 24] and an image prompt ${\mathbf{c}}$. From here, we apply style guidance (Equation 10) with the same prompt ${\mathbf{c}}$ and new style prompt ${\mathbf{c}}_{\text{style}}$ for the desired style application.
Real Image Editing.   Our method can also stylize real images. We use DDIM inversion [87, 24] to invert images into their noisy latent codes at some intermediate step using a reference prompt ${\mathbf{c}}$. From here, we use our style guidance (Equation 10) with reference prompt ${\mathbf{c}}$ and new prompt ${\mathbf{c}}_{\text{style}}$ to denoise the noisy latent code to a stylized image. In Figure 9, we show real image editing results. We provide more details in Appendix C.
@@
Figure 10: Blending multiple style guidances. We can compose multiple customized models by directly blending each style guidance together. Adjusting the blending strength of each model allows us to acquire a smooth style transition. Each stylized image corresponds to different style guidance strengths. Train Image Pair 0 style image credits: Jack Parkhouse
@@
Figure 11: Limitations. Top: our method can cause structure changes in some instances, like change of body position or background changes. Bottom: Our method can change the content in some cases from pretrained output, like the addition of facial hair. We display Baseline DB LoRA for comparison.
@@ 5 Discussion and Limitations
Limitations.   First, our method may occasionally fail to completely maintain input structure, as demonstrated in Figure 11. This could occur as background/pose change (Top), or as additional features being added (Bottom).
@@
Acknowledgments.   We would like to thank Ali Jahanian, Gaurav Parmar, Ruihan Gao, Sean Liu, and Or Patashnik for their insightful feedback and input that contributed to the finished work. We also thank Jack Parkhouse and Aaron Hertzmann for providing style images. Maxwell Jones is supported by the Rales Fellowship and the Siebel Scholar fellowship. This project is partly supported by the Amazon Faculty Research Award, NSF IIS-2239076, Open Philanthropy, and the Packard Fellowship.
@@
In Section A, we evaluate our method against baselines on the diversity metric, showing that our method leads to more diverse generations comparatively. We also show more qualitative results along with a comparison to the concurrent work of Style Aligned Image Generation [31]. In Section B, we then present details of our style guidance formulation. In Section C, we provide implementation details for real image editing, as well as quantitative and qualitative comparisons of real image editing with our method vs baselines. Finally, in Section D, we provide more implementation details, including the setup for our human preference study and the full synthetic training dataset used for evaluation.
Appendix A More Quantitative and Qualitative Results
Style Guidance Ablation   We compare our method, DB LoRA [36, 77], and sliders [23] with both style guidance and LoRA scale used at inference. We also compare against IP Adapters [100] without T2I adapter conditioning, and our Styledrop [86] implementation. In Figure 12, style guidance outperforms the LoRA scale for Ours (blue vs orange), DB LoRA (green vs. pink), and Concept Sliders (brown vs. purple) while our final method still pareto dominates all others, highlighting the effectiveness of both our training scheme and our inference scheme. In Figure 13, our method is still preferred greater than 60 percent of the time against all baselines, while baselines that use style guidance for inference have a higher success rate than those using DB LoRA.
Diversity metric.   To measure the overfitting behavior of our method and baselines, we consider a diversity metric. Concretely, we measure the DreamSim [18] perceptual distance between any two images trained with the same style image pair and generated with the same prompt. We then average results over training pairs and prompts. More formally, we let
DreamSim Diversity (12)
$=$ $\mathbb{E}_{S\in\mathcal{S},P\in\mathcal{P}}\left[\mathbb{E}_{i_{1},i_{2}\in\text{data}_{S,P}}\text{DreamSim}(i_{1},i_{2})\right]$
@@ Appendix A More Quantitative and Qualitative Results
Figure 12: Quantitative comparison with baselines on learned style. Given a fixed inference path, our method’s pareto dominates baselines for image generation both on the same category as training (left) and when evaluated on categories different from training, e.g., trained on human portraits but tested on dog images (right). Secondly, our proposed style guidance outperforms standard LoRA weight scale guidance for our training method (blue vs. orange), DB LoRA (green vs. pink), and Sliders (brown vs. purple). Increased marker size corresponds to an increase in guidance scale.
Figure 13: Human preference study. Our method is preferred over the baselines ($\geq 60\%$), including those using our style guidance. The Gray dashed line denotes 50% chance performance.
@@
Style Aligned Image Generation [31] Baseline   Style Aligned Image Generation [31] is a recent work for zero-shot style-consistent image generation from an exemplar style image. Given the exemplar style image, it is first inverted to a noise map; then for a new text prompt, the image is generated by attending to both its own self-attention map and the self-attention map from the style exemplar at every denoising step. We compare against this baseline by using the style image in our training image pair as an exemplar and generating a new style image with a new text prompt using this method. Optionally, we condition this generation on the edge map of the newly generated image without attention sharing using ControlNet [102] to help with content preservation. We show the qualitative results of our method compared to all the variants of this baseline in Figure 6 of the main paper. Figures 16 and 17 show quantitative comparison, where our method outperforms this baseline in terms of both style similarity and diversity metric. We achieve lower perceptual distance to the style ground-truth images, low perceptual distance from content images, and high diversity.
Extra Qualitative Evaluation   We compare our method with the highest-performing baselines, but use our style guidance (Equation 10 of the main paper ) to apply stylization during inference for these baselines. We present our results in Figure 14. First, we notice that using style guidance for adding style allows the baseline methods to better preserve original content over LoRA scale (Figure 14 vs Figure 5 in the main paper ). While adding our style guidance is better able to preserve content while applying style for baseline methods, our full method is still able to outperform baselines with style guidance applied.
Appendix B Style Guidance Details
@@
$P({\mathbf{x}}|c_{t},c_{\text{style}})$ (13)
$=$ $\frac{P({\mathbf{x}},c_{t},c_{\text{style}})}{P(c_{t},c_{\text{style}})}$
$=$ $\frac{P(c_{\text{style}}|c_{t},{\mathbf{x}})P(c_{t}|{\mathbf{x}})P({\mathbf{x}})}{P(c_{t},c_{\text{style}})}$
@@
$=$ $\ \nabla_{\mathbf{x}}\log\left(\frac{P(c_{\text{style}},c_{t},{\mathbf{x}})}{P(c_{t},{\mathbf{x}})}\right)$
$+\nabla_{\mathbf{x}}\log\left(\frac{P(c_{t},{\mathbf{x}})}{P({\mathbf{x}})}\right)$
$+\nabla_{\mathbf{x}}\log\left(P({\mathbf{x}})\right)$
@@
where $c_{t}$ is the original text prompt, ${\mathbf{c}}_{\text{style}}$ is the class of stylized images from the training style, $\theta_{\text{style}}$ is the UNet with style LoRA adapters applied, and $c_{t,\text{style}}=$$\{{\mathbf{c}}_{t}\}$ in ¡desc¿ style”. Here, we use $c_{t}$ to push the prediction in the text direction, and both text conditioning (“in ¡desc¿ style”) and low-rank adapters ($\theta_{\text{style}}$) to push the prediction into the class of images in the artist’s style denoted by $c_{\text{style}}$. Following this, our new score estimate is:
$\hat{\epsilon}_{\theta}$ $({\mathbf{x}}_{t},c_{t},{\mathbf{c}}_{\text{style}})$
$=$ $\epsilon_{\theta}({\mathbf{x}}_{t},\varnothing)$ (17)
@@
$\lambda_{\text{cfg}}$ and $\lambda_{\text{style}}$ are guidance scales that can be varied as in classifier-free guidance [34]. Given a fixed $\lambda_{\text{cfg}}$, we can vary the $\lambda_{\text{style}}$ term as desired to generate an original guidance $\lambda_{\text{cfg}}$ image with varying amounts of style. Notice that at $\lambda_{\text{cfg}}=\lambda_{\text{style}}$, the $\epsilon_{\theta}({\mathbf{x}}_{t},\varnothing)$ terms cancel and we are left with the original classifier-freeguidance.
Appendix C Real Image Editing Details
Figure 18: Quantitative comparison with baselines on real image editing. Given a fixed classifier-free guidance scale, our method pareto dominates baselines for image generation. During inference, we use style guidance for our method, classifier free guidance for DB LoRA and Sliders, and Image Guidance for IP Adapter w/ T2I. Increased marker size corresponds to an increase in guidance scale.
Figure 19: Real image editing comparison. We compare our real image editing results to other baselines. Our method is able to best preserve content from the real image while applying style, across both training style pairs and input images. We set classifier-free guidance to 5 in all cases. For our method, we set our style guidance (Eq. 10 in the main body) to 6. For baselines, we set LoRA scale (Eq. 3 in the main body) to 1. We invert all images using DDIM inversion [87, 24].
We provide a quantitative and qualitative analysis of our real image editing results compared to baselines. We provide quantitative results in Figure 18. For quantitative evaluation, we use a set of 17 real images from the B-LoRA dataset [17] and add style with all trained style models (see Figure 20 for all training pairs). We measure the perceptual distance to content and style images with DreamSim [18]. Our method Pareto dominates other baselines, yielding both lower perceptual distance to style images and lower perceptual distance to the original content image. We provide qualitative results against baselines in Figure 19. Our method is able to better preserve the real image structure while applying the style from the training pair. We invert the noise to 600 steps for our method and 700 steps for baselines, as we find experimentally that baselines do not apply style when real images are only inverted to 600 steps. We use ReNoise Inversion [24] for our DDIM inversion implementation on SDXL.
Appendix D Implementation Details
@@
We describe the specific methods used to create the paired dataset in detail.
@@
Style descriptor Ablation   We consider replacing a text style descriptor (i.e., ”digital art style”) with a random rare token (i.e., ”$S*$ style”). Applying the same style strength, we quantitatively find that using words yields better distance to content and GT style image ($0.340\pm 0.03$, $0.194\pm 0.03$) than using rare tokens ($0.348\pm 0.04$, $0.209\pm 0.03$), averaged over all test cases. These experiments confirm the remark in StyleDrop [86] that using a text style descriptor produces better results.
@@
Table 1: Experiment Hyperparameters. We choose a fixed stylization hyperparameter for our own model and each baseline when generating images for Mechanical Turk. When picking a hyperparameter, we try and optimize tradeoffs between style application and content preservation, informed by Figure 12. Our style guidance (Equation 10 in the main body ) generally takes values from 0 to $\lambda_{\text{cfg}}=5$, while all other stylization hyperparameters generally take values 0 to 1.